OpenStudy (anonymous):

Divide. x^2 + 2x + 1 / x -2 / x^2 - 1 / x^2 - 4 A. (x + 1) (x + 2) / x - 1 B. (x - 1) (x - 2) / x + 1 C. (x + 1) (x - 2) / x - 1

OpenStudy (pfenn1):

\[\left( x^2+2x+1\over x-2 \right) \over \left( x^2-1 \over x^2 -4 \right)\]Is this the equation?

OpenStudy (pfenn1):

\[\left[ \left( x^2+2x+1\over x-2 \right) \over \left( x^2-1 \over x^2 -4 \right) \right]=\]\[\left( x^2+2x+1\over x-2 \right)\left( x^2-4 \over x^2 -1\right)=\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Now you need to factor all the terms.

OpenStudy (pfenn1):

what are the factors for \[x^2 + 2x +1\]

OpenStudy (pfenn1):

I'll give you a hint:The factors are in the form of (x + a)(x + b) where ab = 1 and a + b = 2